1.) Mark drank a total of 20 liters of water over 4 days. How many days will it take Joe to consume 35 liters of water?
2.) Mom made 72 truffles in 8 days. How many days will it take her to make 99 truffles?

3.) Molly used 22 centimeters of tape to wrap 11 presents. How much tape will she need in all if she has to wrap 12 presents?
4.) The teacher bought 10 books during 5 days of the book fair. After 9 days, how many total books will the teacher have bought?

Answer by ortarico(23)

-------------------
1. 4/20 = x/35
20x = 140
x = 7 i.e. 7days
------------------
2. 8/72 = a/99
72a = 792
a = 11 i.e. 11days
-------------------
3. 11/22 = 12/y
11y = 264
y = 24 i.e. 24cm
------------------
4. 5/10 = 9/z
5z = 90
z = 18 i.e. 18books